Poof:

2. The table shows several animals that live in or below a tree. The table also shows the elevation of their dwellings, relative to ground level. Dwellings Elevation (ft) Robin's nest 12 Groundhog burrow -2 Red squirrel den 3 (a) Plot and label the elevation of each dwelling on a vertical number line. (b) Calculate thePl vertical distance between each pair of dwellings. Organize your results in a table. Answer:

Poof:

Please Help me

Ultrilliam:

@Vocaloid

Poof:

Well this is depressing ;-;

563blackghost:

Hrmmm \(\bf{a.}\). So you need to write this on a number line that is vertical. The highest elevation is 12 while the lowest is -2, so you would label a line from -5 to 15. Labeling the graph would be just like labeling a horizontal graph.  \(\bf{b.}\) Now you need to determine the distances between each. You would subtract the distance between `Robin and Groundhog`, `Robin and and squirrel`, and `Groundhog and squirrel`. So to do that you follow as: Robin and Groundhog distance: \(\bf{12-(-2)=}\) Robin and Squirrel: \(\bf{12-3=}\) Groundhog and Squirrel: \(\bf{-2-3=}\)

Poof:

Robin and Groundhog distance: 12−(−2)= Robin and Squirrel: 12−3= Groundhog and Squirrel: −2−3=

Poof:

how?

Poof:

Whats the = sign for?

Poof:

HEEEEELP